# CLIC: Contrastive Learning Framework for Unsupervised Image Complexity Representation

### v1

@@ Abstract
As an essential visual attribute, image complexity affects human image comprehension and directly influences the performance of computer vision tasks. However, accurately assessing and quantifying image complexity faces significant challenges. Previous works needed more generalization capabilities and well-labeled datasets to learn image complexity features. However, creating such datasets requires expensive manual labeling costs, and the models inevitably learn about human subjective biases. To address the above problems, we propose CLIC, an unsupervised framework based on contrastive learning, for learning image complexity representations. The method learns image complexity features on unlabeled data, avoiding the high labeling cost. Specifically, we propose a unique positive and negative sample selection strategy to reinforce the differences in complexity features. At the same time, we introduce an image prior-based Complexity-Aware Loss to constrain the learning process of the model. We conducted extensive experiments for verification, and the results show that CLIC can effectively learn the image complexity representation. CLIC obtained competitive results with supervised methods by fine-tuning on IC9600. In addition, CLIC applied to downstream tasks shows significant performance improvements, demonstrating the potential for application in various real-world scenarios. code
@@ 1 Introduction
Figure 1: CLIC framework consists of a query encoder and a key encoder with the same structure, and parameter updates are performed by gradient and momentum, respectively. Images in the mini-batch are pre-computed with global entropy as a prior information to construct an auxiliary task. We extract the feature activation energy of the last down-sampling or the last stage with the global entropy of the image to compute the Complexity-Aware loss. In this way, the model is guided to learn the basic features of the image instead of the category or object attributes.
As the complexity of visual tasks increases, accurately assessing and understanding the complexity of images plays a crucial role in downstream task performance (see Appendix A.2). Image Complexity (IC) is a description of the complexity of an image, which is originally a relatively abstract concept that humans can understand [13]. From a subjective point of view, an IC image’s complexity is reflected in its diversity of visual elements, structural complexity, and information density, which together determine the difficulty of human cognition and understanding of the image [24]. Objectively, IC can be regarded as a variety of diverse features in an image, such as color change, texture distribution, and detail representation [44]. Therefore, IC is also one of the crucial attributes to focus on when processing images based on image data-driven computer vision (CV). In recent years, researchers have proposed the concept of computable image complexity to simulate and quantify human visual perception of image complexity through computer algorithms. The quantization and automatic prediction of image complexity have essential roles and applications in several fields, such as image understanding [43], object detection [15], UI design [46], text detection [17], and image enhancement [30].
In order to assess or predict image complexity, some of the earliest work focused on heuristic-based metrics such as information entropy [49], image compression ratio [31], and machine learning methods such as SVMs, random forests, and BP neural networks [9, 19]. However, these algorithms are mainly focus on hand-crafted features, hindering their ability to evaluate the generalization of ICs in practice [12]. Recently, deep convolutional neural networks (CNNs) have demonstrated strong representation and generalization capabilities to explicitly model human subjective perception in a data-driven manner (e.g., image quality evaluation [45], etc.). Several large-scale datasets for evaluating ICs and deep-learning-based IC prediction methods have emerged (e.g., the ComplexityNet [28], ICNet [12], ICCORN [20], etc.). However, building large-scale supervised IC evaluation datasets requires a vast labor cost because of the need to give the ground truth of IC for each image manually. Despite the significant progress made by models such as ICNet, the reliance on manual scoring, with huge labeling costs and subjectivity bias, limits their diffusion in practical applications.
To avoid expensive labeling costs and subjective bias, a very intuitive idea is to utilize unlabeled datasets for IC representation learning, after which the IC score of each image can be computed by fine-tuning with a small number of samples. Self-supervised learning (SSL) [33] generates pseudo-labels by utilizing unlabeled data or creating auxiliary tasks to extract helpful information from the internal structure of the data, which is suitable for situations where labeled data is scarce. SSL can learn valuable features from unlabeled data and reduce the dependence on labeled data. MoCo [23] used momentum encoders and dynamic dictionaries that solve the negative sample insufficiency problem by maximizing the positive sample similarity and minimizing the positive and negative sample similarity through contrastive learning. The success of MoCo triggered extensive subsequent research, such as MoCo v2 [4], SimCLR [3], BYOL [16], and SwAV [1], which pushed forward the development of self-supervised contrastive learning and demonstrated great potential in the field of computer vision.
Self-supervised learning effectively solves the problem of representation learning under unlabeled conditions by constructing pseudo-labels from the data itself. However, extending this approach to the domain of image complexity assessment still faces challenges because of the abstraction and uniqueness of IC features. Driven by the above issues, we propose the contrastive learning framework CLIC (Fig.1) to efficiently learn image complexity features without relying on labeling. Following [23], we construct a query encoder and key encoder, and update the parameters by gradient and momentum. We design a unique positive and negative sample selection strategy for the image complexity representation task. We introduce image prior information to construct a complexity-aware loss, which excludes category interference and guides the model to learn image complexity features accurately. We conducted extensive experiments for validation, and the experimental results show that CLIC can effectively learn the image complexity representation. CLIC fine-tuned in IC9600 yields competitive results with supervised learning. In addition, CLIC applied to downstream tasks shows significant performance improvements, demonstrating the potential for application in various real-world scenarios.
In summary, we make the following contributions:
@@
  • CLIC is proposed as an unsupervised learning framework for image complexity representation, designed to learn the complexity features of images without relying on manual scoring labels.
  • We design a reasonable positive and negative sample selection strategy for the characteristics of image complexity, taking the view closest to the original IC as the positive sample to enhance the contrastive learning for the complexity representation.
  • We propose a complexity-aware loss based on the image, which is a prior information, to optimize the model’s complexity-aware ability by penalizing the difference with the image.
  • Extensive experiments are conducted to demonstrate that CLIC effectively learns image complexity representations. Additionally, CLIC is applied to downstream tasks, resulting in improved performance in both detection and segmentation.
2 Related Works
2.1 Image Complexity Evaluation
Traditional Manual Method. Chipman et al. [7] investigated the complexity factors of visual patterns at a very early stage, and the structural features of similarity transformations and color transformations were able to reduce the perceptual complexity. Ichikawa et al. [26] and Oliva et al. [38] investigated the intrinsic processes underlying human judgments of pattern complexity, exploring how real-world visual complexity is perceived by the human perceptual system. Rosenholtz et al. [41] proposed three measures for the visual clutter problem, and visual clutter is also one of the forms of image complexity. Ramanarayanan et al. [40] used multidimensional scaling MDS to analyze the image complexity. Dimensional scaling MDS to analyze subjects’ judgments of scene complexity. Chikhman et al. [6] used image spatial features, frequency features, and other feature analysis methods to build a computational model to estimate image complexity and compare it with the results given by the observers. Palumbo et al. [39] and Machado et al. [36] both quantified visual complexity based on Image compression algorithms to quantify visual complexity and correlation analysis was used to assess the difference between the quantified results and the subjective complexity.
Intelligent Learning Method. Nagle et al. [37] used a pre-trained convolutional neural network (CNN) to predict complexity ratings for 4000 images, ultimately achieving results that were highly correlated with human observer ratings. Saree et al. [42] explored the extraction of unsupervised information from the intermediate convolutional layers of a CNN, proposing an activation energy metric to quantify visual complexity. Kyle-Davidson et al. [29] used a region neighborhood graph cutting algorithm to capture visually salient regions in an image, computed the local symmetry of an image to extract symmetry features, and built a ComplexityNet to predict image complexity. Feng et al. [12] spent a high cost of labeling to construct an image complexity assessment system that contains 9,600 well-labeled image complexity assessment datasets and provided a baseline model ICNet for predicting image complexity scores and estimating complexity heat maps in a weakly supervised manner.
2.2 Contrastive Learning
InstDisc [48], one of the early approaches to contrast learning, achieves initial representation learning by unsupervisedly treating each image instance as a unique class. MoCo [23] uses momentum encoders to construct a dynamic negative sample pool to solve the problem of insufficient negative samples in mini-batch training. MoCo stores and updates negative samples through a momentum update queue, making the comparison learning perform well in mini-batch training. SimCLR [3] generates the same image with different viewpoints as a pair of positive samples through data augmentation and utilizes large-scale batch processing and projection head networks to enhance the learning effect. Its innovation lies in the data enhancement strategy and large-scale training. SwAV [1] performs comparative learning utilizing clustering, utilizing image clustering results of different perspectives as the comparison target and using online clustering as the comparison target. Results as comparison targets, combining online clustering with comparison learning, enabling the model to perform more efficient representation learning unsupervised. BYOL [16] achieves self-guided learning of representations by alternating the training of two networks (online and target networks). No explicit negative sample pairs are required, demonstrating high-quality representation learning. DINO [50] transfers the knowledge from the teacher’s network to the student’s network using a self-supervised distillation method to achieve efficiency without labeling learning.
@@ 3.1 Positive and Negative Samples Selection
A judicious selection of positive and negative samples in representation learning can help the model learn meaningful image representations. This meaning is generally defined for a specific task (e.g., similarity of pairs of positive samples for image classification). Therefore, in IC representation learning, we propose a positive and negative samples selection strategy for IC representations to ensure that the model can focus on the complex features of images rather than the content, categories, or object properties. Unlike traditional representation learning, positive and negative sample selection for IC representations needs to be designed around the complexity properties of the images themselves, rather than the similarity of categories or content, or even try to exclude these interfering factors as much as possible.
In the image complexity assessment task, each image needs to be assigned a score to represent its image complexity. Theoretically, the IC scores of any two images will not be equal. Therefore, we consider each image a separate class, i.e., images other than themselves are other classes and are considered negative samples. Some work, such as SimCLR [3], generates different views from the original image as positive sample pairs and images outside the mini-batch as negative samples. We also follow this approach (Fig.2(a)). Nevertheless, the difference is that we do special treatment for the views.
@@
Figure 2: (a) Positive and Negative Samples Selection. We crop the mini-batch image to produce views whose image complexity is close ($\sim$ IC score) to that of the original as positive samples. Outside the mini-batch are negative samples. (b) Small-scale crop and merge. Given an image of size $(h,w)$, the crop size is $(h,w)/c$, $(h,w)/2c$, and $(h,w)/4c$, the corresponding number of crops for each image is $c$, $2c$, and $4c$, respectively. Then, we randomly merge two crops of the same size and their two transformed crops.
@@
Positive sample pairs for the image classification task belong to one class, i.e., the class of randomly crop sub-images from the image is the same class as that of the original image. Moreover, the IC scores of the original image and the resulting view should be as close as possible in IC representation. Therefore, we design crop schemes with different sizes (with data augmentation) and evaluate them with the Pearson correlation coefficient (PCC) [18] between the views and the original image. Finally, the crops with the closest PCC is used as the views and defined as a positive sample pair.
@@
Our framework has no additional definition for negative samples and considers images outside the mini-batch as negative samples while building a dynamic queue to store negative sample features [23]. In addition, we process the original dataset by small-scale crop and merge it in such a way as to introduce local image features for the model learning phase, which will be more beneficial for IC representation learning [12]. Note that small-scale crop and merge (Fig.2(b)) is not just a boost to the original data volume but is more focused on local feature enhancement, allowing the model to learn a localized complexity representation of the image.
@@ 3.2 Complexity-Aware Loss with Global Entropy Prior
Contrastive loss, commonly used in representation learning, can distinguish similarities or differences between positive and negative samples. In contrast, similarity computation tends to introduce image category attributes for model learning, which may result in complexity perception errors. Therefore, image complexity representation learning needs to guide the model to focus more on or even prioritize the primary attributes of an image (which should be independent of category, content, and object) when extracting features. We found an essential phenomenon with some simple metrics for image complexity assessment (Tab.1). We randomly sampled 1000 images from the ImageNet training set, computed the IC evaluation results of each simple metric for these samples separately, and used the inference output of ICNet as the ground truth. We found that edge density, UAE [42], and global entropy all have relatively high Pearson correlation coefficients with the ground truth (details in Appendix C), and earlier, some scholars have similarly shown that metrics such as global entropy can evaluate image complexity well [36].
@@
In addition, Saraee et al. [42] investigated the feature extraction properties of intermediate convolutional layers utilizing deep neural networks capable of capturing both low-level (e.g., edges, texture) and high-level features (e.g., objects and scene content) of an image. They quantize several convolutional feature maps to create a mapping of feature activation energy to image complexity scores in an unsupervised form. Surprisingly, our experiments found the highest PCC between UAE and global entropy, compared to other combinations, and a highly positive correlation with ground truth. A relationship between global entropy and feature activation energy can be established to guide the model in learning a representation close to the ground truth.
Driven by these observations, we consider using an image’s global entropy ge as a prior information to construct an auxiliary task for image complexity representation learning, proposing the image Complexity-Aware Loss (CAL). The goal is to optimize the complexity properties of positive and negative samples. The image complexity characterization task needs to focus the model on image complexity features and use the image prior information to make the model generate representations that minimize the content, category, or object attributes prone to complexity-awareness errors. For each image in the mini-batch, its global entropy is pre-computed using the following form:
@@
Where $N$ is the batch size, $i\in[0,N]$. $x_{i}^{fae}$ and $x_{i}^{ge}$ denote feature activation energy fae and global entropy ge, respectively. In addition, following some representation learning methods [23, 4, 3], this paper also considers InfoNCE [47] loss form ${L_{IN}}$ (Eq.(4)) to measure positive and negative sample similarity in terms of dot product while building dictionary look-up problem for contrastive learning.

### v2

@@ Abstract
As a fundamental visual attribute, image complexity significantly influences both human perception and the performance of computer vision models. However, accurately assessing and quantifying image complexity remains a challenging task. (1) Traditional metrics such as information entropy and compression ratio often yield coarse and unreliable estimates. (2) Data-driven methods require expensive manual annotations and are inevitably affected by human subjective biases. To address these issues, we propose CLIC, an unsupervised framework based on Contrastive Learning for learning Image Complexity representations. CLIC learns complexity-aware features from unlabeled data, thereby eliminating the need for costly labeling. Specifically, we design a novel positive and negative sample selection strategy to enhance the discrimination of complexity features. Additionally, we introduce a complexity-aware loss function guided by image priors to further constrain the learning process. Extensive experiments validate the effectiveness of CLIC in capturing image complexity. When fine-tuned with a small number of labeled samples from IC9600, CLIC achieves performance competitive with supervised methods. Moreover, applying CLIC to downstream tasks consistently improves performance. Notably, both the pretraining and application processes of CLIC are free from subjective bias.
keywords:
Representation learning , contrastive learning , image complexity , image prior
journal: Pattern Recognition
[…]
@@ 1 Introduction
Image Complexity (IC) describes the level of visual complexity present in an image—a concept that is inherently abstract but intuitively understandable by humans [1]. Subjectively, IC is reflected in the diversity of visual elements, structural intricacy, and information density within an image, all of which influence the cognitive load required for human interpretation [2]. Objectively, IC can be characterized by various visual features, including color variations, texture distributions, and detail richness [3]. Given its intrinsic connection to visual perception, IC has emerged as a critical attribute in numerous computer vision (CV) tasks. In recent years, the notion of computable image complexity has been introduced, aiming to algorithmically simulate and quantify human perception of image complexity. Accurate quantification and automatic prediction of IC have shown significant potential in a range of applications, such as image understanding [4], object detection [5], UI design [6], text detection [7], and image enhancement [8]. Therefore, accurate assessment and modeling of image complexity are essential not only for aligning with human perception but also for enhancing the performance of downstream CV tasks (Sec. 3.2). A reliable IC estimation facilitates more effective and perceptually aligned processing in various computer vision pipelines.
@@
Figure 1: The CLIC framework comprises a query encoder and a key encoder that share the same architecture. The parameters of the query encoder are updated via standard backpropagation, while the key encoder is updated using momentum-based updates. Each image in a mini-batch is pre-computed with its global entropy (ge) as prior information. To guide the model towards learning intrinsic image features rather than category- or object-specific attributes, we extract the feature activation energy (fae) from the last down-sampling layer (or the final stage of the encoder). This fae is then combined with the image’s ge to compute the Complexity-Aware Loss. This loss encourages the model to focus on structural and informational content that reflects image complexity, rather than semantic content.
To assess or predict image complexity, early research primarily focused on heuristic-based metrics such as information entropy [9], image compression ratio [10], and classical machine learning techniques including SVMs, random forests, and BP neural networks [11, 12]. However, these methods largely rely on hand-crafted features, which limit their capacity to generalize image complexity (IC) assessment in real-world scenarios [13]. In recent years, deep convolutional neural networks (CNNs) have shown strong representational power and generalization ability, enabling more effective modeling of human subjective perception in a data-driven manner—particularly in related tasks such as image quality assessment [14]. This progress has led to the development of large-scale IC evaluation datasets (e.g., IC9600 [13]) and corresponding deep learning-based IC prediction models (e.g., ComplexityNet [15], ICNet [13], ICCORN [16], etc.). Nevertheless, constructing such large-scale supervised datasets remains labor-intensive, as IC scores must be manually annotated for each image, posing a major bottleneck in the scalability and practicality of supervised IC learning approaches.
To mitigate the high cost and inherent subjectivity associated with manual annotations, a natural alternative inspired by unsupervised learning is to leverage unlabeled datasets for image complexity (IC) representation learning. Once meaningful representations are learned, IC scores can be obtained by fine-tuning the model on a small set of labeled samples. In the realm of unsupervised learning, contrastive learning methods—such as MoCo [17]—employ momentum encoders and dynamic dictionaries to address the challenge of insufficient negative samples. These methods aim to maximize the similarity between positive pairs while minimizing the similarity between positive and negative pairs. The success of MoCo has catalyzed the development of numerous enhanced approaches, including MoCo v2 [18], SimCLR [19], BYOL [20], and SwAV [21]. These methods have significantly advanced the field of self-supervised contrastive learning and demonstrated strong performance across a wide range of computer vision tasks.
While contrastive learning has shown great success in unsupervised representation learning, directly applying it to image complexity (IC) representation poses significant challenges. (1) IC features are abstract and distinct from general image features; they should remain invariant across different image categories and not be dominated by category-specific attributes. (2) Defining positive and negative pairs for IC representation is non-trivial, as conventional contrastive learning on classification datasets (e.g., ImageNet) tends to induce category bias into the learned representations. Motivated by these challenges, we propose the CLIC framework (Fig. 1), which leverages contrastive learning to effectively learn image complexity representations in an unsupervised manner. Specifically, CLIC employs a dual-encoder structure consisting of a query encoder and a key encoder, updated via gradient descent and momentum, respectively. To address the difficulty of sample pair construction in the context of IC, we design a novel positive and negative sample selection strategy tailored for complexity representation learning. Moreover, we incorporate image prior information to formulate a Complexity-Aware Loss, which explicitly suppresses category-related interference and guides the model to focus on structural and informational cues associated with complexity. Extensive experiments validate the effectiveness of the proposed framework. CLIC, when fine-tuned on a limited number of labeled samples from IC9600 [13], achieves performance competitive with fully supervised methods. Furthermore, applying CLIC to various downstream tasks yields significant performance improvements, highlighting its practical value and broad applicability in real-world scenarios.
In summary, our main contributions are as follows:
  • 1.
@@
    CLIC, a novel contrastive learning framework for unsupervised image complexity representation, is proposed. This framework enables the learning of complexity-related features without relying on manual scoring labels.
  • 2.
    A tailored positive and negative sample selection strategy is designed, specifically for image complexity. The augmented view that most closely preserves the original image’s complexity is treated as the positive sample for contrastive loss calculation.
  • 3.
    A Complexity-Aware Loss, based on global entropy as prior knowledge, is introduced. This loss function optimizes the model’s ability to capture complexity by penalizing discrepancies between the learned representation and the image’s inherent complexity.
  • 4.
    Extensive experiments demonstrate the effectiveness of CLIC. The simple weighted application of CLIC encoder features significantly improves performance across various downstream computer vision tasks.
2 Related Work
Image Complexity Estimation. Modeling image complexity has long been a topic of interest in both computer vision and human perception. Early works relied on handcrafted features, such as entropy [22], symmetry [23, 24], spatial layout [25], and compressibility [26, 27], to quantify complexity in a rule-based manner. These methods provide interpretable signals but struggle to generalize to diverse real-world images.
With the advent of deep learning, learning-based approaches have emerged to model visual complexity. Nagle et al. [28] and Saree et al. [29] utilized CNN features to regress complexity scores obtained from human annotations. ComplexityNet [24] leveraged local symmetry descriptors, while Feng et al. [13] proposed ICNet, a weakly supervised model trained on a curated complexity dataset (IC9600). Extensions of ICNet have incorporated ordinal regression [16] or segmentation statistics [30, 31, 32] to better capture structural aspects. Despite their effectiveness, these models require either human supervision or proxy labels, limiting their scalability and adaptability to new tasks.
Contrastive Representation Learning. Contrastive learning has become a standard paradigm for self-supervised visual representation learning. Methods such as InstDisc [33], MoCo [17], and SimCLR [19] learn representations by pulling together augmented views of the same instance while pushing apart different instances. Further refinements like BYOL [20], SimSiam [34], and SwAV [21] improve training stability and efficiency by removing the need for negative samples or incorporating clustering. These methods have demonstrated strong performance in downstream recognition tasks.
However, contrastive learning has primarily been explored in the context of semantic representation learning, where the objective is to encode discriminative features for downstream classification or detection. Existing frameworks are not designed to capture structural complexity, which is often orthogonal to semantic content.
Our Perspective. In this work, we explore contrastive learning from a new perspective—learning to represent visual complexity in an unsupervised manner. To the best of our knowledge, this is the first attempt to adapt contrastive learning for image complexity modeling. Our proposed CLIC framework introduces a complexity-aware sampling strategy and a complexity-aligned loss function, guiding the encoder to capture structural and perceptual difficulty rather than semantic categories. This enables learning complexity representations without reliance on labeled data, offering a scalable and task-agnostic solution for perceptual modeling.
3 Preliminary
3.1 Definition of Image Complexity
Image complexity is inherently difficult to define. Traditional manual methods, such as image entropy and image compression ratio, offer some metrics, but these are primarily heuristic and often inaccurate for assessing complexity, with limited generalization ability. Machine learning methods typically rely on supervised learning to model image complexity, which requires costly annotations and inevitably incorporates human subjective biases. Regardless of whether manual or machine learning methods are employed, image complexity modeling can be formally expressed as in Eq. (1), where ${x_{i}}$ represents an image from a dataset and $f(\cdot)$ denotes the mapping from this image to its corresponding image complexity $IC_{i}$, with $IC_{i}$ taking values within the range of 0 to 1.
$I{C_{i}}=f({x_{i}})$ (1)
After analyzing the image complexity across several datasets, it is observed that, in general, the Image Complexity Distribution (ICD) of a dataset approximates a normal distribution, i.e., $IC_{i}\sim N(\mu,\sigma^{2})$, suggesting that most images tend to exhibit a medium level of complexity. As shown in Fig. 2, the ICD of the dataset is computed using both image entropy and machine learning methods, and the resulting distribution curve closely follows the normal distribution pattern. Notably, datasets with a normal characteristic ICD are predominantly composed of natural scene images, such as street scenes, indoor/outdoor environments, and fields. However, certain datasets, like the Cornell Grasp Dataset [35], exhibit a high concentration of ICD values around 0.17. This characteristic is one of the reasons why tasks involving such datasets, particularly the 2D target detection task on the Cornell Grasp Dataset, tend to be easier to learn.
Figure 2: Image Complexity Distribution of Datasets. (a) global entropy of MS COCO [36]. (b) ICD of MS COCO [36]. (c) ICD of PASCAL VOC [37]. (d) ICD of Cornell Grasp [35]. The global entropy is obtained by Eq.(2). ICD stands for image complexity distribution, which represents the true complexity of each image. In this work, it is obtained by conducting statistics on the inference of each image by a well-trained ICNet.
IC Range AP AP50 AP75
All 40.5 59.3 43.8
<0.3 41.2(+0.7) 60.3(+1.0) 44.6(+0.8)
>0.7 39.6(-0.9) 58.0(-1.3) 42.6(-1.2)
Table 1: Results of YOLOX-S on MS COCO with different IC range. The training details for YOLOX-S were maintained consistent with the official implementation. All represents the original MS COCO val set, with the IC range from 0 to 1.
3.2 Impaction of Image Complexity
[…]
@@ 4.1 Positive and Negative Samples Selection
Figure 3: Our overall architecture. (a) Detail structure of CLIC. (b) Positive and Negative Samples Selection. We crop the mini-batch image to produce views whose image complexity is close ($\sim$ IC score) to that of the original as positive samples. Outside the mini-batch are negative samples.
A careful selection of positive and negative samples in representation learning plays a critical role in helping the model learn meaningful image representations. This meaning is typically defined for a specific task, such as the similarity of positive sample pairs in image classification. In the context of image complexity (IC) representation learning, we propose a strategy for selecting positive and negative samples that ensures the model focuses on the complexity features of images, rather than on content, categories, or object properties. Unlike traditional representation learning, the selection of positive and negative samples for IC representation must be designed around the intrinsic complexity properties of the images themselves, rather than relying on category or content similarity. In fact, it is crucial to minimize interference from such extraneous factors as much as possible.
In the image complexity assessment task, each image is assigned an IC score to represent its complexity. Theoretically, the IC scores of any two images are unlikely to be identical. As a result, we treat each image as belonging to a distinct class, meaning that all images except for the image itself are considered negative samples. Some methods, such as SimCLR [19], generate different views of the original image as positive sample pairs and treat images outside the mini-batch as negative samples. We adopt a similar approach (Fig. 3(b)), with a key distinction being that we apply special treatment to the views in the context of IC representation learning.
Figure 4: Small-scale crop and merge. Given an image of size $(h,w)$, the crop size is $(h,w)/c$, $(h,w)/2c$, and $(h,w)/4c$, the corresponding number of crops for each image is $c$, $2c$, and $4c$, respectively. Then, we randomly merge two crops of the same size and their two transformed crops.
@@
In the image classification task, positive sample pairs belong to the same class, meaning that sub-images obtained through random cropping from the original image are considered part of the same class as the original image. Additionally, the IC scores of the original image and the resulting view should be as similar as possible in the IC representation. To achieve this, we design cropping schemes of varying sizes (with data augmentation) and evaluate them using the Pearson correlation coefficient (PCC) [39] between the views and the original image. The crops with the highest PCC are selected as the views and are defined as a positive sample pair. It is important to note that this IC score proximity refers to the fact that, from a statistical perspective, we have designed specific combinations to obtain detailed cropping schemes. As a result, cropping is directly applied during training, eliminating the need for additional computational time overhead.
@@
Our framework (Fig.3(a)) does not introduce any additional definition for negative samples; instead, images outside the mini-batch are treated as negative samples, with a dynamic queue used to store negative sample features [17]. Furthermore, we process the original dataset by applying small-scale cropping and merging it in a way that introduces local image features into the model learning phase. This process enhances the learning of image complexity representations [13]. It should be noted that the small-scale crop and merge (Fig.4) approach does not simply augment the original data volume but focuses on enhancing local features, allowing the model to learn a more localized complexity representation of the image.
@@ 4.2 Complexity-Aware Loss with Global Entropy Prior
Contrastive loss, commonly used in representation learning, is designed to distinguish similarities or differences between positive and negative samples. However, similarity computation in such methods may inadvertently introduce image category attributes into the model’s learning process, leading to potential errors in complexity perception. Therefore, image complexity representation learning must guide the model to focus more on, or even prioritize, the fundamental attributes of an image—attributes that should be independent of category, content, or object—when extracting features. We observed an important phenomenon using some simple metrics for image complexity assessment (Tab. 2). Specifically, we randomly sampled 1000 images from the ImageNet training set, computed the IC evaluation results for each of these simple metrics, and used the inference output of ICNet as the ground truth. The results showed that edge density, UAE [29], and global entropy all exhibit relatively high Pearson correlation coefficients with the ground truth (details in B). This finding is consistent with prior work, where scholars have demonstrated that metrics like global entropy can effectively evaluate image complexity [27].
@@
In addition, Saraee et al. [29] explored the feature extraction properties of intermediate convolutional layers in deep neural networks, which are capable of capturing both low-level features (e.g., edges, texture) and high-level features (e.g., objects and scene content) of an image. They quantized several convolutional feature maps to create a mapping from feature activation energy to image complexity scores in an unsupervised manner. Interestingly, our experiments revealed that the highest Pearson correlation coefficient (PCC) was found between UAE and global entropy, compared to other combinations, and that there was a certain degree of positive correlation with the ground truth. This suggests that a relationship between global entropy and feature activation energy can be established to guide the model in learning a representation that aligns closely with the ground truth.
Motivated by these findings, we propose using an image’s global entropy ge as prior information to construct an auxiliary task for image complexity representation learning. This leads to the formulation of the Image Complexity-Aware Loss (CAL). The objective of CAL is to optimize the complexity characteristics of both positive and negative samples. The image complexity characterization task should direct the model’s focus toward the image complexity features, utilizing the image’s prior information to generate representations that minimize errors caused by content, category, or object attributes—errors that could hinder the model’s complexity awareness. For each image in the mini-batch, its global entropy is pre-computed using the following formula:
@@
where $N$ is the batch size, $i\in[0,N]$. $x_{i}^{fae}$ and $x_{i}^{ge}$ denote feature activation energy fae and global entropy ge, respectively.
In addition, following some representation learning methods [17, 18, 19], this paper also considers InfoNCE [42] loss form ${L_{IN}}$ (Eq.(5)) to measure positive and negative sample similarity in terms of dot product while building dictionary look-up problem for contrastive learning.